# BMGQ: A Bottom-up Method for Generating Complex Multi-hop Reasoning Questions from Semi-structured Data

### v1

@@ 3.2.2 Processing Pipeline
The Node Information Construction stage (Figure 1 Part 2) builds a structured, semantically coherent neighborhood around each seed entity to support subsequent evidence chain expansion. This process involves lightweight text preprocessing followed by a core filtering step based on named entity recognition (NER).
@@ 3.3 Evidence Chain Construction
(b) Evidence graph generated from the same seed (“Japan Airlines”) after incorporating NLI-based relation classification, logical diversity, and semantic diversity constraints.
@@ 3.3.2 Graph Expansion: Controlled Breadth-First Growth
Figure 1 Part 3 provides a schematic overview of this pipeline, showing how relation classification guarantees the validity of edges, diversity evaluation prevents redundancy, and breadth-first expansion yields a balanced and interpretable evidence cluster.
The algorithm operates layer by layer, starting from a seed node and expanding outward according to a user-specified strategy (e.g., $[4,2,2]$, meaning four nodes at depth 1, two per node at depth 2, and so on). At each step, three main components govern the expansion:
@@
For example, starting from the seed entity Japan Airlines, the expansion yields a multi-layered graph that links not only to other airlines, but also to related people (e.g., Shingo Katori), organizations (e.g., SMAP), locations (e.g., Tokyo, Sapporo), and cultural events (e.g., Kōhaku Uta Gassen). In addition, the system enriches the seed with discriminative attributes—such as founding year (1951), hub (Osaka–Kansai), and alliance membership (Oneworld)—so that these attribute cues alone can uniquely identify the correct seed entity (Japan Airlines). This demonstrates that the system avoids semantic homogeneity, instead producing logically coherent and semantically diverse chains that serve as high-quality raw material for downstream multi-hop question generation.
3.4 Final Question Construction
The final stage of our pipeline transforms the evidence clusters generated in Step 3 into complex multi-hop questions. The input consists of the evidence cluster (nodes, edges, and supporting passages) and the seed entity (the answer). Unlike naive question generation, which often yields factoid-style queries easily solved by direct retrieval, our method employs a bottom-up, reverse reasoning strategy. This ensures that the constructed question faithfully reproduces the reasoning path encoded in the evidence cluster.
@@ 3.4.1 Reverse Question Generation
We adopt a bottom-up approach that begins from the leaf nodes of the evidence cluster and progressively backtracks to the seed node. At each step, the model is prompted to convert local evidence into a descriptive but oblique clue, which avoids explicitly mentioning the seed or its immediate neighbors. These clues are then recursively nested to form a composite reasoning chain, culminating in a single question that uniquely points to the seed.
@@ 3.4.3 Iterative Refinement
This rewriting procedure is grounded in the underlying evidence graph. First, the system selects a minimal supporting subgraph containing multiple deep cues (i.e., nodes located several hops away from the seed, beyond the most immediate neighbors). To prevent shallow retrieval, anchor terms and “killer pairs” (e.g., award + year, model + maker) are softened or replaced with implicit formulations. Rather than directly naming entities, the system rephrases them using higher-level categories, relations, and event descriptors, ensuring that no direct level-1 node or alias appears in the surface text.
@@
Through this loop, the discriminative properties required to uniquely identify the seed are preserved, but the search complexity is progressively increased. This ensures that even as the question becomes harder, the seed remains the only correct answer. The loop terminates once the question reaches the target difficulty level or the maximum number of refinement rounds.
@@ 4.2.1 Prediction-Based Validation
In the first stage, the constructed question is answered by multiple inference runs, yielding a set of candidate predictions $\{\text{seed},\text{prediction1},\text{prediction2},\dots\}$. If the majority of predictions match the seed answer, the question is considered stable and accepted directly. Otherwise, it is flagged for further analysis, together with the seed and all alternative predictions. This serves as a coarse filtering mechanism, effectively removing questions with unclear or inconsistent reasoning.
@@ 4.2.2 Constraint Decomposition and Evidence-Based Verification
    • If only the seed remains, the question is accepted at this stage.
    • If multiple candidates remain, they are ranked based on their explicit-match score $S_{\text{exp}}$. If the seed achieves the uniquely highest score, the question is accepted; otherwise, it moves to fine-grained verification.
@@
    • If the seed achieves the uniquely highest $S_{\text{norm}}$, the question is retained; otherwise, it is discarded.
@@ 5 Conclusion
Our approach introduces a four-stage pipeline encompassing data adaptation, node construction, evidence chain generation, and question synthesis, transforming raw Wikipedia dumps into logically rich, semantically diverse, and structurally coherent reasoning graphs. Furthermore, we propose a Data Quality Evaluation System that integrates multi-model consensus, structured clue decomposition, and evidence-based verification. This system ensures that only questions with unique answers, consistent logic, and verifiable evidence are preserved, significantly improving the reliability and discriminative power of the generated dataset.
By automating the construction of BrowseComp-level multi-hop datasets, our framework reduces the prohibitive cost of manual curation while enabling scalable production of challenging, high-quality training data. This advancement has the potential to accelerate research on reasoning-centric large language models and facilitate the development of models capable of complex, multi-step inference.

### v2

@@ 3.2.2 Processing Pipeline
The Node Information Construction stage (the part 2 of the Figure 1) builds a structured, semantically coherent neighborhood around each seed entity to support subsequent evidence chain expansion. This process involves lightweight text preprocessing followed by a core filtering step based on named entity recognition (NER).
@@ 3.3 Evidence Chain Construction
(b) Evidence graph generated from the same seed entity(“Japan Airlines”) after incorporating NLI-based relation classification, logical diversity, and semantic diversity constraints.
@@ 3.3.2 Graph Expansion: Controlled Breadth-First Growth
The part 3 of the Figure 1 provides a schematic overview of this pipeline, showing how relation classification guarantees the validity of edges, diversity evaluation prevents redundancy, and breadth-first expansion yields a balanced and interpretable evidence cluster.
The algorithm operates layer by layer, starting from a seed entity node and expanding outward according to a user-specified strategy (e.g., $[4,2,2]$, meaning four nodes at depth 1, two per node at depth 2, and so on). At each step, three main components govern the expansion:
@@
For example, starting from the seed entity Japan Airlines, the expansion yields a multi-layered graph that links not only to other airlines, but also to related people (e.g., Shingo Katori), organizations (e.g., SMAP), locations (e.g., Tokyo, Sapporo), and cultural events (e.g., Kōhaku Uta Gassen). In addition, the system enriches the seed entity with discriminative attributes—such as founding year (1951), hub (Osaka–Kansai), and alliance membership (Oneworld)—so that these attribute cues alone can uniquely identify the correct seed entity (Japan Airlines). This demonstrates that the system avoids semantic homogeneity, instead producing logically coherent and semantically diverse chains that serve as high-quality raw material for downstream multi-hop question generation.
3.4 Question Construction & Optimization
The final stage of our pipeline transforms the evidence clusters generated in Step 3 into complex multi-hop questions. The input consists of the evidence cluster (nodes, edges, and supporting passages) and the seed entity. Unlike naive question generation, which often yields factoid-style queries easily solved by direct retrieval, our method employs a bottom-up, reverse reasoning strategy. This ensures that the constructed question faithfully reproduces the reasoning path encoded in the evidence cluster.
@@ 3.4.1 Reverse Question Generation
We adopt a bottom-up approach that begins from the leaf nodes of the evidence cluster and progressively backtracks to the seed entity node. At each step, the model is prompted to convert local evidence into a descriptive but oblique clue, which avoids explicitly mentioning the seed entity or its immediate neighbors. These clues are then recursively nested to form a composite reasoning chain, culminating in a single question that uniquely points to the seed entity.
@@ 3.4.3 Iterative Refinement
This rewriting procedure is grounded in the underlying evidence graph. First, the system selects a minimal supporting subgraph containing multiple deep cues (i.e., nodes located several hops away from the seed entity, beyond the most immediate neighbors). To prevent shallow retrieval, anchor terms and “killer pairs” (e.g., award + year, model + maker) are softened or replaced with implicit formulations. Rather than directly naming entities, the system rephrases them using higher-level categories, relations, and event descriptors, ensuring that no direct level-1 node or alias appears in the surface text.
@@
Through this loop, the discriminative properties required to uniquely identify the seed entity are preserved, but the search complexity is progressively increased. This ensures that even as the question becomes harder, the seed entity remains the only correct answer. The loop terminates once the question reaches the target difficulty level or the maximum number of refinement rounds.
@@ 4.2 Data Quality Evaluation Workflow
Figure 4: A two-step data quality evaluation pipeline. The evaluation process consists of: (1) Multi-Model Answer Validation, and (2) Structured Clue Decomposition & Evidence-Based Verification.
@@ 4.2.1 Prediction-Based Validation
In the first stage, the constructed question is answered by multiple inference runs, yielding a set of candidate predictions $\{\text{seed entity},\text{prediction1},\text{prediction2},\dots\}$. If the majority of predictions match the seed entity, the question is considered stable and accepted directly. Otherwise, it is flagged for further analysis, together with the seed entity and all alternative predictions. This serves as a coarse filtering mechanism, effectively removing questions with unclear or inconsistent reasoning.
@@ 4.2.2 Constraint Decomposition and Evidence-Based Verification
    • If only the seed entity remains, the question is accepted at this stage.
    • If multiple candidates remain, they are ranked based on their explicit-match score $S_{\text{exp}}$. If the seed entity achieves the uniquely highest score, the question is accepted; otherwise, it moves to fine-grained verification.
@@
    • If the seed entity achieves the uniquely highest $S_{\text{norm}}$, the question is retained; otherwise, it is discarded.
5 Experiments
To evaluate both the difficulty and answerability of our automatically constructed multi-hop questions, we conducted a two-part experimental study. The first part measures model performance across several representative datasets to assess whether our generated questions achieve comparable reasoning depth to real-world retrieval-intensive benchmarks. The second part performs a graph-based structural analysis to quantify the reasoning and relational complexity of each dataset through linguistic and topological metrics.
5.1 Model Evaluation on Multi-Hop Datasets
We evaluated multiple large language models – GPT-5 (Advanced Reasoning), Gemini-2.5-Pro, Doubao-1.6-Thinking, and Deepseek-V3.2 – on four datasets: HotpotQA, ComplexWebQuestions, BrowseComp, and our BMGQ dataset. Each model independently generated answers, and correctness was measured by strict exact-match comparison against reference answers. The overall results are summarized in Table 2:
Table 2: Model performance comparison across datasets.
Dataset / Model GPT-5 Gemini-2.5-Pro Doubao-1.6-Thinking Deepseek-V3.2
HotpotQA 76% 49% 43% 39%
ComplexWebQuestions 55% 56% 50% 33%
BrowseComp 23% 12% 1% 1%
BMGQ Dataset 37% 13% 5% 3%
Across all models, accuracy drops sharply from HotpotQA and ComplexWebQuestions to BMGQ dataset and BrowseComp, revealing a substantial increase in reasoning and retrieval difficulty. While HotpotQA primarily involves shallow two-hop reasoning, both BMGQ and BrowseComp dataset emphasize hard-to-search but easy-to-verify questions – where individual clues remain ambiguous, but their combination uniquely constrains the answer space.
GPT-5 maintains the highest overall accuracy, yet its performance on our BMGQ dataset (37%) remains close to that on BrowseComp (23%), confirming that our automatically generated questions approach evaluation-level difficulty. Meanwhile, models such as Gemini-2.5-Pro and Doubao-1.6-Thinking show particularly steep performance declines, underscoring that these datasets demand multi-domain reasoning and longer retrieval chains, which exceed the capacity of models primarily tuned for short-hop factual QA.
Overall, the results demonstrate that our automatic construction framework successfully produces questions comparable in complexity to leading evaluation datasets while retaining scalability for training use.
5.2 Graph-based Structural Analysis
To characterize the structural and semantic complexity of different multi-hop QA datasets, we construct a graph representation for each question, as described in Section 4.1. Each question is converted into a directed heterogeneous graph (G = (V, E)), where nodes represent objects (entities) or attributes (properties, events, conditions), and edges represent one of six linguistically grounded relation types. We compute a set of graph-theoretic and semantic metrics over these graphs to quantify dataset difficulty along the axes of breadth, depth, constraint richness, and relational complexity. Below, we detail each metric, its computation, and the capability dimension it captures.
(1) Objects (#).
  • Definition. Number of object nodes extracted from the prompt.
  • Interpretation. Measures the breadth and depth of the reasoning space: more objects imply that the question touches more entities, domains, or facts, increasing the search space and making shortcut retrieval less viable.
  • How it is obtained. Directly counted from the graph’s object-type nodes.
(2) Attributes (#).
  • Definition. Number of attribute nodes representing properties, states, temporal descriptors, or contextual modifiers.
  • Interpretation. Reflects constraint richness: more attributes impose stricter filtering conditions and help enforce answer uniqueness.
  • How it is obtained. Count the attribute-type nodes extracted during relation parsing.
(3) Number of Edges.
  • Definition. Total number of undirected relations among nodes in (E).
  • Interpretation. Higher edge density typically indicates more interdependencies between clues, which increases reasoning entanglement and reduces the risk of shallow single-hop resolution.
(4) Average Degree.
  • Definition. $\text{AvgDegree}(G)=\frac{1}{|V|}\sum_{v\in V}\deg(v)$
  • Interpretation. Captures the overall interconnectedness of the semantic structure. A higher value suggests that nodes participate in multiple relations, increasing combinatorial complexity.
  • How it is computed. Degree is counted in an undirected sense to reflect relational connectivity regardless of linguistic directionality.
(5) Maximum Degree.
  • Definition. $\text{MaxDegree}(G)=\max_{v\in V}\deg(v)$
  • Interpretation. Indicates the presence of hub nodes, often representing pivotal clues that connect multiple reasoning branches. Datasets with higher max-degree tend to involve more intricate, non-linear reasoning.
(6) Graph Diameter.
  • Definition. Longest shortest-path distance between any pair of nodes in the undirected graph.
  • Interpretation. Measures the depth of reasoning: larger diameters correspond to longer chains of dependencies and more multi-step inference.
(7) Longest Path Starting from the Subject.
  • Definition. Length of the longest simple path originating from the designated subject node.
  • Interpretation. Captures targeted depth: how far a solver must propagate from the main subject to integrate peripheral clues.
  • How it differs from diameter. Diameter measures global structure; this metric isolates subject-centered reasoning chains.
(8) Number of Linear Logical Paths.
  • Definition. Count of simple, non-branching paths in which every intermediate node has degree 2.
  • Interpretation. Represents the number of independent reasoning chains embedded in a question. More paths → broader search space and more compositional subproblems.
(9) Number of Cycles.
  • Definition. Number of simple cycles in the undirected graph.
  • Interpretation. Cycles reflect interdependent reasoning, where clues reinforce each other; datasets with more cycles tend to resist shortcut single-hop strategies.
(10) Number of Intersection Nodes.
  • Definition. Nodes with $\text{degree}\geq 3$ serving as junctions connecting multiple reasoning threads.
  • Interpretation. A high count indicates branching complexity: solvers must integrate multiple paths rather than follow one linear chain.
(11) Mean Relation-Type Diversity.
  • Definition. Average number of distinct relation types (among the six defined in Section 4.1) per question: $\text{RelationDiversity}=\frac{1}{6}\sum_{r}\text{count}(r)$
  • Interpretation. Quantifies the balance and distributional evenness of relation types across a question’s evidence graph—datasets with more uniformly distributed causal, taxonomic, compositional, attributive, temporal/locational, and other relations avoid over-reliance on any single relation category and better reflect structurally varied reasoning patterns.
(12) Undirected Degree Centrality (Score).
  • Definition. $C_{D}(v)=\frac{\deg(v)}{|V|-1}$, aggregated across nodes (mean).
  • Interpretation. Indicates how influence is distributed across the graph; high centrality suggests concentrated evidence hubs requiring prioritization during reasoning.
(13) Undirected Betweenness Centrality (Score).
  • Definition. $C_{B}(v)=\sum_{s\neq v\neq t}\frac{\sigma_{st}(v)}{\sigma_{st}}$, where ($\sigma_{st}$) is the number of shortest paths between (s) and (t).
  • Interpretation. Measures whether reasoning passes through narrow “choke points” — high scores imply pivotal clues that cannot be bypassed.
(14) Undirected Eigenvector Centrality (Score).
  • Definition. Centrality proportional to the centrality of neighbors: $x_{i}=\frac{1}{\lambda}\sum_{j}A_{ij}\,x_{j}$.
  • Interpretation. Captures global importance: clues connected to other influential clues receive high values. Datasets with elevated eigenvector centrality often encode deeply nested or hierarchical reasoning.
To evaluate whether our proposed BMGQ dataset achieves the structural and inference complexity characteristic of challenging multi-hop benchmarks, we further compare its graphical representation with three representative datasets: ComplexWebQuestions, HotpotQA, and BrowseComp. Table 3 lists the averages of fourteen structural and semantic metrics. The results show a clear progression in complexity from earlier multi-hop benchmarks to BrowseComp, and then to our BMGQ dataset.
Metric ComplexWebQuestions HotpotQA BrowseComp BMGQ
Objects 1.25 1.69 6.69 10.20
Attributes 0.77 1.80 6.09 6.07
Number of Edges 2.94 4.56 14.69 17.83
Average Degree 1.47 1.64 1.96 1.95
Maximum Degree 2.58 3.28 6.03 6.35
Graph Diameter 2.29 2.98 5.42 6.84
Longest Path Starting from the Subject 1.91 2.58 5.37 5.62
Number of Linear Logical Paths 1.58 2.20 5.16 6.92
Number of Cycles 0.03 0.11 0.88 0.67
Number of Intersection Nodes 0 0.03 0.42 0.31
Mean Relation-Type Diversity 0.49 0.76 2.47 2.95
Undirected Degree Centrality (Score) 0.92 0.77 0.46 0.39
Betweenness Centrality (Score) 0.93 0.82 0.73 0.76
Eigenvector Centrality (Score) 0.68 0.64 0.59 0.59
Table 3: Graph-structural metrics.
Breadth of reasoning.
BMGQ dataset exhibits the largest number of object nodes (10.20) among all datasets, nearly doubling that of BrowseComp (6.69) and far exceeding HotpotQA and ComplexWebQuestions (<2<2). This indicates that BMGQ questions integrate a substantially wider set of entities, covering more domains and forcing the solver to explore a broader search space. The edge count follows a similar trend: BMGQ has 17.83 edges on average—significantly larger than BrowseComp (14.69) and an order of magnitude higher than early benchmarks (<5<5). This demonstrates that BMGQ questions encode more relational interactions and richer factual interdependencies.
Depth and multi-hop reasoning complexity.
Both graph diameter and longest path starting from the subject quantify the depth of reasoning chains. BMGQ shows the highest values (diameter = 6.35; longest subject path = 5.62), closely mirroring BrowseComp (6.03 and 5.37, respectively) and far surpassing HotpotQA or ComplexWebQuestions ($\approx 3$). This reflects that BMGQ, like BrowseComp, requires solvers to traverse long, multi-stage reasoning chains rather than relying on shallow or local evidence.
Branching and compositional complexity.
The average degree (1.95) and maximum degree (6.84) of BMGQ again align closely with BrowseComp (1.96 and 6.03), indicating comparably dense local connectivity. Meanwhile, the numbers of linear logical paths and intersection nodes capture branching structure: BMGQ shows 6.92 linear chains and 0.31 intersection nodes, whereas BrowseComp shows 5.16 and 0.42, respectively. These values together indicate that BMGQ questions encompass multiple parallel reasoning threads with moderate merging complexity – comparable to BrowseComp yet more structured than early datasets, which exhibit minimal branching ($\approx 0-0.1$ intersection nodes).
Cyclic and interdependent reasoning.
Cycles reinforce mutual constraints and reduce shortcut-solving. BrowseComp contains 0.88 cycles per question, while BMGQ has 0.67, substantially higher than HotpotQA or ComplexWebQuestions ($\leq 0.11$). This shows that BMGQ maintains interlocking logical structures rather than simple linear chains, a hallmark of difficult retrieval-based reasoning tasks.
Semantic richness and relation diversity.
An important feature of real-world multi-hop reasoning is the diversity of relational semantics. BrowseComp shows an average of 2.47 relation types per question, while BMGQ reaches an even higher diversity of 2.95. In contrast, early datasets employ almost exclusively taxonomic or compositional relations (<0.8<0.8). This demonstrates that BMGQ questions integrate a wide mix of attributive, causal, compositional, temporal, and classificatory relations, creating heterogeneous reasoning requirements similar to or exceeding those in BrowseComp.
Centrality-based difficulty.
The centrality metrics—degree, betweenness, and eigenvector centrality – characterize how crucial particular nodes are in mediating the flow of reasoning.
  • In all three centrality scores, BMGQ (0.76/0.76/0.59) nearly matches BrowseComp (0.73/0.46/0.59).
  • BMGQ’s slightly higher degree and betweenness scores indicate more influential hubs and more mandatory intermediate reasoning steps than in BrowseComp.
  • Early benchmarks again show inflated degree centrality (due to small graph size) but low betweenness and eigenvector centrality, reflecting shallow and loosely connected structures.
Overall assessment.
Across all fourteen metrics, BMGQ consistently clusters alongside BrowseComp and diverges sharply from early multi-hop datasets. BMGQ matches or exceeds BrowseComp in:
  • reasoning breadth (objects, edges),
  • depth (diameter, longest path),
  • relation diversity, and
  • centrality-driven structural difficulty.
Meanwhile, BrowseComp retains slightly higher cyclic and intersection-node complexity, which is expected given its focus on evaluation-only problem design. Nonetheless, the overall structural fingerprints of the two datasets are strongly aligned.
5.3 Human Quality Assessment
While model performance and graph-structural analyses jointly demonstrate that our generated questions resemble high-difficulty benchmarks in complexity and reasoning depth, it remains essential to verify the intrinsic quality of the constructed data itself. To this end, we conduct a human evaluation to assess whether the generated questions are well-formed, logically sound, and paired with correct and uniquely identifiable answers.
We randomly sampled a substantial subset of questions from the full dataset and asked human annotators to evaluate them along three dimensions:
  1. 1.
    Question soundness: whether the question is well-posed, free of contradictions, and contains no malformed or misleading clues.
  2. 2.
    Answer correctness: whether the provided seed answer is factually correct and supported by evidence.
  3. 3.
    Answer uniqueness: whether the clues in the question uniquely identify the seed answer, without allowing alternative plausible candidates.
Annotators evaluated each example by examining the natural-language question together with all explicit constraints stated in the prompt. Their task was to determine whether these constraints, taken collectively, uniquely support the labeled answer (i.e., the seed entity). An item was marked as a failure only if at least one of the following conditions occurred: the question contained contradictory or incoherent cues; the seed entity could not be verified as satisfying all stated conditions; or multiple candidate entities were found to satisfy every constraint in the prompt, indicating lack of answer uniqueness.
Across the sampled subset, the dataset achieved an overall pass rate of 87%, indicating that the vast majority of automatically constructed questions are logically coherent, correctly labeled, and uniquely solvable. This level of quality is particularly notable given the difficulty of the target task and the minimal amount of human intervention involved. The remaining 13% of cases typically involved edge scenarios—such as clue over-abstraction or insufficiently restrictive combinations—which can be further alleviated by tightening the constraints in the question optimization stage or adjusting the thresholds in the quality evaluation pipeline.
Together with the strong model-based difficulty evaluation and the alignment of graph-structural metrics with BrowseComp-level complexity, the human assessment confirms that our generated dataset is not only challenging but also reliable, verifiable, and high-quality, validating the effectiveness of the BMGQ framework.
@@ 6 Conclusion
To this end, we introduced BMGQ, a Bottom-up Method for Generating Complex Multi-hop Reasoning Questions from semi-structured data. BMGQ comprises a four-stage pipeline encompassing data adaptation, node construction, evidence chain generation, and question synthesis, transforming semi-structured knowledge sources into logically rich, semantically diverse, and structurally coherent reasoning graphs. Furthermore, we propose a Data Quality Evaluation System that integrates multi-model consensus, structured clue decomposition, and evidence-based verification. This system ensures that only questions with unique answers, consistent logic, and verifiable evidence are preserved, significantly improving the reliability and discriminative power of the generated dataset.
By automating the construction of BrowseComp-level multi-hop datasets, BMGQ reduces the prohibitive cost of manual curation while enabling scalable production of challenging, high-quality training data. This advancement has the potential to accelerate research on reasoning-centric large language models and facilitate the development of models capable of complex, multi-step inference.